# SpeakerLM: End-to-End Versatile Speaker Diarization and Recognition with Multimodal Large Language Models

### v2

@@ Abstract
The Speaker Diarization and Recognition (SDR) task aims to predict “who spoke when and what” within an audio clip, which is a crucial task in various real-world multi-speaker scenarios such as meeting transcription and dialogue systems. Existing SDR systems typically adopt a cascaded framework, combining multiple modules such as speaker diarization (SD) and automatic speech recognition (ASR). The cascaded systems suffer from several limitations, such as error propagation, difficulty in handling overlapping speech, and lack of joint optimization for exploring the synergy between SD and ASR tasks. To address these limitations, we introduce SpeakerLM, a unified multimodal large language model for SDR that jointly performs SD and ASR in an end-to-end manner. Moreover, to facilitate diverse real-world scenarios, we incorporate a flexible speaker registration mechanism into SpeakerLM, enabling SDR under different speaker registration settings. SpeakerLM is progressively developed with a multi-stage training strategy on large-scale real data. Extensive experiments show that SpeakerLM demonstrates strong data scaling capability and generalizability, outperforming state-of-the-art cascaded baselines on both in-domain and out-of-domain public SDR benchmarks. Furthermore, experimental results show that the proposed speaker registration mechanism effectively ensures robust SDR performance of SpeakerLM across diverse speaker registration conditions and varying numbers of registered speakers [^1] Demo page: https://sites.google.com/view/speakerlm/overview.
@@ 1 Introduction
  • Extensive empirical validations and analysis. To evaluate the effectiveness of SpeakerLM, we build various cascaded SDR baselines with open-source SD and ASR models that achieve SOTA performance on different benchmarks. Results show that SpeakerLM demonstrates strong data scaling capability, outperforming all the baselines on both in-domain and out-of-domain public benchmarks.
@@ 2.2 Speaker Diarization and Recognition
Figure 1 compares three system architectures for SDR. Figure 1 (a) depicts a conventional SDR system that pipelines SD and ASR modules (sd+asr1_2024), denoted as SD+ASR. This pipeline has several limitations, such as error propagation and lack of joint optimization. To address these issues, one direction uses LLMs for post-processing (sd_llm_2_2025), and another direction jointly models SD and ASR tasks within a single unified model in an end-to-end manner. We denote these two methods by SD+ASR+LLM and E2E-SDR, respectively.
@@ 5.1 Implementation Details
We use the AdamW (adamw_2024) optimizer for optimization. The learning rate is initially set to 1e-5, linearly increased to 5e-5 during the warm-up phase, and then gradually decayed following a cosine schedule. We use a dynamic batching strategy, adjusting the number of samples per batch based on their length, with a maximum token limit of 6K to balance efficiency and memory usage. Models are trained on 4 NVIDIA A800 GPUs for 1M steps at each stage, with a validation step of 10K. The training pseudocode and the specific prompts used for LLMs are presented in demo page.
@@ 5.3 Performance without Speaker Registration
Initializing SpeakerLM by SpeakerLM-ASR, we conduct the following training with different scales of SDR data, ranging from 212.25 hours (Alimeeting-Train+AISHELL4-Train) to 7,638.95 hours (Alimeeting-Train+AISHELL4-Train+In-House-Train). Results show that when trained on a limited amount of SDR data, SpeakerLM falls behind most cascaded baselines. However, as the training data scale up, SpeakerLM demonstrates a strong data scaling capability, with substantial improvements in cpCER and $\Delta\textrm{cp}$. Moreover, we observe that the improvements on CER are marginal. This is because the in-house data mainly consists of near-field recordings, which offer limited ASR gains for far-field speech with reverberation. With 7,638.95 hours training data, SpeakerLM significantly outperforms all baselines. In terms of cpCER, SpeakerLM achieves absolute improvements of 6.60%, 6.56%, and 13.82% over the strongest cascaded system on AliMeeting-Eval, AISHELL-4-Eval, and AISHELL-5-Eval, respectively. Notably, SpeakerLM also achieves superior performance on the out-of-domain and challenging test set, AISHELL5-Eval, with a $\Delta\textrm{cp}$ of 0.57, indicating strong robustness and generalizability of SpeakerLM to unseen and noisy acoustic environments.

### v3

@@ Abstract
The Speaker Diarization and Recognition (SDR) task aims to predict “who spoke when and what” within an audio clip, which is a crucial task in various real-world multi-speaker scenarios such as meeting transcription and dialogue systems. Existing SDR systems typically adopt a cascaded framework, combining multiple modules such as speaker diarization (SD) and automatic speech recognition (ASR). The cascaded systems suffer from several limitations, such as error propagation, difficulty in handling overlapping speech, and lack of joint optimization for exploring the synergy between SD and ASR tasks. To address these limitations, we introduce SpeakerLM, a unified multimodal large language model for SDR that jointly performs SD and ASR in an end-to-end manner. Moreover, to facilitate diverse real-world scenarios, we incorporate a flexible speaker registration mechanism into SpeakerLM, enabling SDR under different speaker registration settings. SpeakerLM is progressively developed with a multi-stage training strategy on large-scale real data. Extensive experiments show that SpeakerLM demonstrates strong data scaling capability and generalizability, outperforming state-of-the-art cascaded baselines on both in-domain and out-of-domain public SDR benchmarks. Furthermore, experimental results show that the proposed speaker registration mechanism effectively ensures robust SDR performance of SpeakerLM across diverse speaker registration conditions and varying numbers of registered speakers.
Codehttps://sites.google.com/view/speakerlm/overview
@@ 1 Introduction
  • Extensive analysis. To evaluate the effectiveness of SpeakerLM, we build various cascaded SDR baselines. Results show that SpeakerLM demonstrates strong data scaling capability, outperforming all the baselines on both in-domain and out-of-domain public benchmarks.
@@ 2.2 Speaker Diarization and Recognition
Figure 1 (a) depicts a conventional SDR system that pipelines SD and ASR modules (sd+asr1_2024), denoted as SD+ASR. This pipeline has several limitations, such as error propagation and lack of joint optimization. To address these issues, one direction uses LLMs for post-processing, and another direction jointly models SD and ASR tasks within a single unified model in an end-to-end manner. As shown in Figure 1, we denote these two methods by SD+ASR+LLM and E2E-SDR, respectively.
@@ 5.1 Implementation Details
We use the AdamW (adamw_2024) optimizer for optimization. The learning rate is initially set to 1e-5, linearly increased to 5e-5 during the warm-up phase, and then gradually decayed following a cosine schedule. We use a dynamic batching strategy, adjusting the number of samples per batch based on their length, with a maximum token limit of 6K to balance efficiency and memory usage. Models are trained on 4 NVIDIA A800 GPUs for 1M steps at each stage, with a validation step of 10K. The training pseudocode and the specific prompts used for LLMs are presented in code page.
@@ 5.3 Performance without Speaker Registration
Initializing SpeakerLM by SpeakerLM-ASR, we conduct the following training with different scales of SDR data, ranging from 212.25 hours (Alimeeting-Train+AISHELL4-Train) to 7,638.95 hours. Results show that when trained on a limited amount of SDR data, SpeakerLM falls behind most cascaded baselines. However, as the training data scale up, SpeakerLM demonstrates a strong data scaling capability, with substantial improvements in cpCER and $\Delta\textrm{cp}$. Moreover, we observe that the improvements on CER are marginal. This is because the in-house data mainly consists of near-field recordings, which offer limited ASR gains for far-field speech with reverberation. With 7,638.95 hours training data, SpeakerLM significantly outperforms all baselines. In terms of cpCER, SpeakerLM achieves absolute improvements of 6.60%, 6.56%, and 13.82% over the strongest cascaded system on AliMeeting-Eval, AISHELL-4-Eval, and AISHELL-5-Eval, respectively. Notably, SpeakerLM also achieves superior performance on the out-of-domain and challenging test set, AISHELL5-Eval, with a $\Delta\textrm{cp}$ of 0.57, indicating strong robustness and generalizability of SpeakerLM to unseen and noisy acoustic environments.